What is the firm's market to book value for 2009?

The formula and answer is as follows:

$100/[($850,000/25,000)] = 2.9400

Please explain where the $100 comes from and why dividing the common stock by total current liabilities. Please provide a thorough explaination of the entire problem.Snapit Company Income Statement (2009) Sales Cost of goods sold Gross profit Selling and administrative expenses Operating profit Interest expense Income before tax Tas expense Net income $4,000,000 3.040,000 960.000 430.000 530.000 160,000 370,000 148,000 $222,000 Balance Sheet 2008 Cash Accounts receivable Inventory Total current assets Fixed assets Total assets 2009 $60,000$50,000 550,000 500,000 690.000 1,300,000 1,170,000 1.300,000 1,230,000 2,600,000 2,400,000 б 20.000 270,000 580,000 500,000 850,000 750,000 900,000 1,000,000 1.750.000 1.750.000 250,000 600.000 $2.600,000 $2.400,000 250,000 Accounts pavable Bank loan Total current liabilities Bonds payable Total liabilities Common stock (25.000 shares) Retained earnings Total liabilities & equity 250,000 400,000 Note: The common shares are trading in the stock market for S100 each
Answer #1

Market ot book value = Market share price / Book value per share

Share is trading at $100 each as mentioned in the question. Therefore market value of per share is $100.

Book value = Total assets - total liabilities

Book value = 2,600,000 - 1,750,000

Book value = $850,000

Book value per share = Book value / total shares outsanding

Book value per share = 850,000 / 25,000

market to book ratio = market value per share / ( book value per share)

market to book ratio = 100 / ( 850,000 / 25,000)

market to book ratio = 2.94